OpenStudy (anonymous):

35=1/3b(5) solve for b

OpenStudy (waleed_imtiaz):

b=35*3/5.. SO u will get... b=21

OpenStudy (anonymous):

35=1/3b(5) solve for b ...wht about this (5) inside the brackets

OpenStudy (anonymous):

the actual formula is v=1/3bh so h is 5

OpenStudy (anonymous):

thank you waleed_imtiaz

OpenStudy (waleed_imtiaz):

my pleasure..... .. :D